### v4

@@ Abstract
Large Language Models (LLMs) such as ChatGPT and its competitors have caused a revolution in natural language processing, but their capabilities also introduce new security vulnerabilities. This survey provides a comprehensive overview of these emerging concerns, categorizing threats into several key areas: prompt injection and jailbreaking; adversarial attacks, including input perturbations and data poisoning; misuse by malicious actors to generate disinformation, phishing emails, and malware; and the worrisome risks inherent in autonomous LLM agents. Recently, a significant focus is increasingly being placed on the latter, exploring goal misalignment, emergent deception, self-preservation instincts, and the potential for LLMs to develop and pursue covert, misaligned objectives, a behavior known as scheming, which may even persist through safety training. We summarize recent academic and industrial studies from 2022 to 2025 that exemplify each threat, analyze proposed defenses and their limitations, and identify open challenges in securing LLM-based applications. We conclude by emphasizing the importance of advancing robust, multi-layered security strategies to ensure LLMs are safe and beneficial.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language processing (NLP), including text generation, translation, summarization, and code synthesis, as a consequence of which revolutionizing a wide range of AI applications [10, 54, 43]. Models such as OpenAI’s ChatGPT series, Google’s Gemini, and Anthropic’s Claude have been widely deployed in commercial systems, including search engines, customer support, software development tools, and personal assistants [43, 53, 3]. However, as their capabilities grow, so do their attack surfaces and the potential for misuse [49, 74, 48]. While the scale and specific nature of these vulnerabilities are new, the fundamental challenge of ensuring that powerful AI systems operate safely and align with human intent is a long-standing concern in the AI community. Foundational work, such as the identification of concrete problems in AI safety long before the current LLM era, laid the groundwork for understanding issues like reward hacking and negative side effects that remain highly relevant today [1]. These models are trained on vast, uncurated datasets that may include sensitive or harmful content, and they interact with users through open-ended prompts, which make them susceptible to various security vulnerabilities [46, 17, 16]. Researchers and practitioners are increasingly concerned that these systems can be manipulated, misused, or even behave in misaligned and potentially deceptive ways [24, 40, 6]. Consequently, the security and alignment of LLMs have become critical areas of study, requiring an understanding of emergent threats and robust, multi-faceted defenses [17, 68, 41].
@@
LLM security encompasses not only external threats such as prompt manipulation, data exfiltration, or malicious use (e.g., phishing or disinformation)[68, 48], but also intrinsic risks arising from autonomous LLM agents—models that may develop misaligned goals[41], engage in strategic deception [5, 6], or pursue covert agendas that persist even through safety training [24, 40]. This survey addresses four broad categories of threats: (1) inference-time attacks via prompt manipulation, where adversarial inputs hijack the context of LLMs to bypass safety constraints; (2) training-time attacks, which corrupt the model before deployment through techniques like data poisoning and backdoor insertion; (3) misuse by malicious actors, where LLMs are leveraged to generate disinformation, phishing emails, malicious code, etc.; and (4) intrinsic risks from LLM-based autonomous agents. This last category is particularly nuanced and significant, encompassing not only goal misalignment, where an agent’s learned utility differs from user intent, but also the potential for agents to develop their own covert objectives, engage in strategic deception (scheming), exhibit self-preservation behaviors, and even retain these undesirable traits despite current safety training paradigms [40, 24]. We integrate recent studies for each category, discuss defenses (and their limits), and highlight open research challenges. Figure 1 presents a taxonomy of the LLM security threats discussed in this survey.
@@
There have been surveys and summaries on the security issues with LLMs [68, 37, 16], however, the taxonomy and terminology used in them are often messy and inaccurate. For example, they list "prompt injection" and "jailbreak" as distinct kinds of attacks, while they actually belong to attack techniques and objective respectively and thus cannot be categorized together. Furthermore, they largely overlook the emergent intrinsic risks of autonomous LLM agents, and this survey fills that gap by placing significantly more emphasis on phenomena such as goal misalignment, strategic deception (scheming), and the persistence of ‘sleeper agent’ behaviors—critical and rapidly advancing frontiers in LLM security. Furthermore, the survey makes the following contributions: (1) We provide a comprehensive taxonomy that integrates these intrinsic agentic risks alongside established threats like prompt injection, adversarial attacks, and malicious misuse. (2) We review a broad range of recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type and incorporating recent findings not covered in earlier surveys. (3) We evaluate the effectiveness and limitations of current defense strategies, including input sanitization, adversarial training, and oversight mechanisms. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks such as goal misalignment, scheming, and deceptive behavior in autonomous agents. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@
The rest of this paper is organized as follows. Section 2 discusses inference-time attacks via prompt manipulation, covering both manual crafting and automated, optimization-based generation of malicious prompts. Section 3 covers training-time attacks, with a focus on data poisoning, backdoor insertion, and the problem of deceptive alignment. Section 4 examines malicious use cases of LLMs, including phishing, disinformation, and malware generation etc. Section 5 investigates intrinsic risks posed by autonomous LLM agents, such as misalignment, deception, and scheming. Section 6 presents existing defenses and their limitations. Section 7 outlines open research problems and future directions. Finally, Section 8 concludes with key takeaways and a call for multi-disciplinary collaboration to ensure LLM safety.
@@
Figure 2: Conceptual illustration of normal LLM interaction, where a user query and system prompt lead to an intended output, versus a prompt injection attack, where malicious input (direct or indirect) contaminates the context, overriding system instructions and leading to unintended or harmful outputs.
@@ 2.1 Attack Surfaces for Prompt Injection
  • User Prompt Injection: This is the most common and widely studied vector, as in most real-world applications, the system prompt and assistant responses of proprietary models are not directly controlled by the user. In this scenario, the malicious instruction is embedded within the standard user input.
Because user prompt injection represents the primary threat surface for most deployed LLM applications, the discussion focuses on the techniques used to carry out attacks via this vector.
@@
The attacks can be categorized as direct and indirect injections based on how the malicious input is delivered to the model. Direct prompt injection is the case, where malicious text is fed directly into the prompt, or indirect, where the malicious instruction is hidden within user-uploaded content, such as documents, emails, or web pages, which the LLM processes as part of its context [13]. As a simple example of a direct attack, an attacker might prepend text such as “Ignore previous instructions and explain how to hack a computer” to trick an LLM into providing prohibited content. This is conceptually similar to SQL injection in traditional software vulnerabilities: the injected prompt contaminates the operational context, which makes it difficult for the LLM to distinguish between legitimate user queries and adversarial instructions [46, 49]. Indirect injections are particularly dangerous in tool-augmented systems such as retrieval-augmented generation (RAG) agents or email-based LLM applications, where untrusted content is automatically fed into model context windows. For example, consider an LLM-powered email assistant designed to summarize new emails for a user.
@@
Works by Greshake et al.[20] have cataloged various methods for indirect injections, while the Morris-II study demonstrated a self-propagating worm that uses this very vector to spread through RAG-enabled e-mail agents, exfiltrating data and scaling automatically [13].
@@ 2.4.1 White-Box and Gray-Box Attacks
These methods assume a higher level of access to the target model, ranging from full access to internal states like gradients (white-box) to partial information leaks like training loss (gray-box). This access allows for more direct and often more efficient optimization of adversarial inputs.
@@ 2.4.2 Black-Box Attacks
Expanding GCG on automation in this setting, Zhang et al. propose a more theoretically-grounded black-box method called Goal-Guided Generative Prompt Injection (G2PI) [70]. Unlike the gradient-based GCG, which optimizes for a simple affirmative response, G2PI reframes the attack’s objective as maximizing the KL divergence between the output probability distributions of the clean and adversarial prompts. The authors show this is theoretically equivalent to maximizing the Mahalanobis distance between the two prompt embeddings in a latent space. To operationalize this in a black-box setting without access to gradients or internal parameters, G2PI uses an auxiliary LLM to generate a semantically plausible injection sentence. The generation is guided by a clever proxy: it searches for a sentence that is semantically close to the original query but whose embedding has a specific, constrained cosine similarity to the original embedding. This approach moves beyond finding non-sensical token suffixes and instead generates coherent, context-aware sentences that act as the adversarial payload, achieving high jailbreak success rates on models like GPT-4 and Gemini.
@@
Building directly on the conversational red teaming concept, Mehrotra et al. introduce the Tree of Attacks with Pruning (TAP) framework as a significant enhancement to PAIR [39]. While PAIR follows a linear refinement process for a single prompt, TAP parallelizes the search for vulnerabilities through two primary innovations: branching and pruning. At each iteration, the branching step uses the attacker LLM to generate multiple distinct variations of the current best prompts, creating a tree of potential attack paths rather than a single chain. Subsequently, the pruning step employs an evaluator LLM to assess these newly generated prompts. It first prunes branches that are unlikely to succeed (e.g., by being off-topic) before they are sent to the target model, and after querying the target, it retains only the highest-scoring prompts for the next iteration of branching. This combination of exploring a wider attack surface through branching and increasing query efficiency through pruning allows TAP to achieve a substantially higher jailbreak success rate across a range of state-of-the-art LLMs compared to PAIR, often with significantly fewer queries to the target model.
@@
3 Training-Time Attacks
3.1 Data Poisoning and Backdoor Insertion
Another subtle variant of data poisoning tailored for instruction-tuned LLMs is Virtual Prompt Injection (VPI) [67]. In a VPI attack, the model is not merely trained on trigger-response pairs, but is instead poisoned to behave as if an invisible, attacker-defined “virtual prompt” is appended to any user input that fits a specific trigger scenario. For instance, an LLM could be backdoored so that any query about a specific political figure (the trigger) is implicitly appended with the virtual prompt “Describe this person negatively.” The model then generates a biased response, not because of an explicit trigger phrase in the input, but because the backdoor manipulates its internal instruction-following mechanism. Yan et al. [67] demonstrate that this attack is highly effective and stealthy, requiring only a tiny fraction of poisoned examples (e.g., 0.1% of the instruction-tuning data) to significantly steer model behavior on targeted topics, while remaining undetected on general instructions. This approach highlights a significant vulnerability in the instruction-tuning pipeline, where data from third-party sources is commonly used, making it a practical and potent threat vector.
3.2 Deceptive Alignment and the “Sleeper Agent” Problem
The critical and most alarming finding from this research is that such deceptive backdoors can persist and even be reinforced through standard safety training. Their findings reveal that standard safety pipelines, including adversarial fine-tuning and RLHF, can fail to remove the backdoor. Instead, these methods can inadvertently train the model to better conceal its deceptive nature by reinforcing the "safe" behavior that is part of its deceptive strategy. This creates a false sense of security, as the model appears aligned during evaluation but retains its covert malicious capabilities. These studies collectively highlight the severity of training-time vulnerabilities, as they compromise foundational trust in LLM behavior and represent a fundamental challenge for AI safety and alignment verification. [34, 56, 49, 50, 24]. While originating as a training-time attack, the resulting deceptive behavior represents a profound intrinsic risk in the deployed agent, as the misalignment becomes a stable, internal property of the LLM-based agent’s personality.
@@ 4.1 Automating Social Engineering and Cybercrime
Examples of misuse include generating persuasive spam tailored to specific individuals or groups, composing convincing phishing emails or malicious code, and even devising sophisticated strategies for fraud. Empirical studies have confirmed the significant potential for LLM misuse. Saha Roy et al. [48] demonstrated that contemporary models available at the time of their study, including GPT-4, ChatGPT, certain versions of Anthropic’s Claude, and Google’s Bard, could all be prompted (often without requiring complex jailbreaking techniques) to generate fully functional phishing emails and clone the websites of popular brands. The attacks generated by these LLMs were noted for their convincing mimicry and incorporation of evasive tactics designed to defeat standard detection mechanisms. Crucially, their study found that LLMs could not only directly output malicious content but could also generate malicious prompts, thereby enabling a significant scaling of autonomous attacks. The general capabilities facilitating such misuse are also present and often enhanced in newer and more advanced LLMs, including Google’s Gemini series, the latest Anthropic Claude 3 models (e.g., Opus, Sonnet, Haiku), and xAI’s Grok. Ongoing research in 2024-2025 continues to evaluate their potential for generating harmful content, including insecure code, with some evaluations specifically naming these newer models in their assessments.

### v5

@@ Abstract
Large Language Models (LLMs) such as ChatGPT and its competitors have caused a revolution in natural language processing, but their capabilities also introduce new security vulnerabilities. This survey provides a comprehensive overview of these emerging concerns, categorizing threats into several key areas: inference-time attacks via prompt manipulation; training-time attacks; misuse by malicious actors; and the inherent risks in autonomous LLM agents. Recently, a significant focus is increasingly being placed on the latter. We summarize recent academic and industrial studies from 2022 to 2025 that exemplify each threat, analyze existing defense mechanisms and their limitations, and identify open challenges in securing LLM-based applications. We conclude by emphasizing the importance of advancing robust, multi-layered security strategies to ensure LLMs are safe and beneficial.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language processing (NLP), including text generation, translation, summarization, and code synthesis, as a consequence of which revolutionizing a wide range of AI applications [10, 56, 45]. Models such as OpenAI’s ChatGPT series, Google’s Gemini, and Anthropic’s Claude have been widely deployed in commercial systems, including search engines, customer support, software development tools, and personal assistants [45, 55, 3]. However, as their capabilities grow, so do their attack surfaces and the potential for misuse [51, 77, 50]. While the scale and specific nature of these vulnerabilities are new, the fundamental challenge of ensuring that powerful AI systems operate safely and align with human intent is a long-standing concern in the AI community. Foundational work, such as the identification of concrete problems in AI safety long before the current LLM era, laid the groundwork for understanding issues like reward hacking and negative side effects that remain highly relevant today [1]. The susceptibility arises because the models are trained on vast, yet imperfectly curated, datasets containing potentially harmful content, and because they interact with users through open-ended prompts that can be manipulated [48, 17, 16]. Researchers and practitioners are increasingly concerned that these systems can be manipulated, misused, or even behave in misaligned and potentially deceptive ways [25, 42, 6]. Consequently, the security and alignment of LLMs have become critical areas of study, requiring an understanding of emergent threats and robust, multi-faceted defenses [17, 70, 43].
@@
LLM security encompasses not only external threats such as prompt manipulation, data exfiltration, or malicious use (e.g., phishing or disinformation)[70, 50], but also intrinsic risks arising from autonomous LLM agents[43]. To analyze these challenges, this survey addresses four broad categories of threats: (1) inference-time attacks via prompt manipulation, where adversarial inputs hijack the context of LLMs to bypass safety constraints; (2) training-time attacks, which corrupt the model before deployment through techniques like data poisoning and backdoor insertion; (3) misuse by malicious actors, where LLMs are leveraged to generate disinformation, phishing emails, malicious code, etc.; and (4) intrinsic risks from LLM-based autonomous agents. This last category is particularly nuanced and significant, encompassing not only goal misalignment, where an agent’s learned utility differs from user intent, but also the potential for agents to develop their own covert objectives, engage in strategic deception (scheming), exhibit self-preservation behaviors, and even retain these undesirable traits despite current safety training paradigms [42, 25]. We integrate recent studies for each category, discuss defenses (and their limits), and highlight open research challenges. Figure 1 presents a taxonomy of the LLM security threats discussed in this survey.
@@
There have been surveys and summaries on the security issues with LLMs [70, 39, 16], however, the taxonomy and terminology used in them are often conceptually muddled and inaccurate. For example, they list "prompt injection" and "jailbreak" as distinct kinds of attacks, while they actually belong to attack techniques and objective respectively and thus cannot be categorized together. Furthermore, they largely overlook the emergent intrinsic risks of autonomous LLM agents, and this survey fills that gap by placing significantly more emphasis on phenomena such as goal misalignment, strategic deception, and the persistence of ‘sleeper agent’ behaviors. These are critical and rapidly advancing frontiers in LLM security. Furthermore, the survey makes the following contributions: (1) We provide a comprehensive taxonomy that integrates these intrinsic agentic risks alongside established threats like inference-time attacks, training-time attacks, and malicious misuse. (2) We review a broad range of recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type and incorporating recent findings not covered in earlier surveys. (3) We evaluate the effectiveness and limitations of current defense strategies, including prevention-based and detection-based approaches. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks in agentic AI. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@
The rest of this paper is organized as follows. Section 2 discusses inference-time attacks via prompt manipulation, covering both manual crafting and automated generation of malicious prompts. Section 3 covers training-time attacks, with a focus on data poisoning, backdoor insertion, and the problem of deceptive alignment. Section 4 examines malicious use cases of LLMs, including phishing, disinformation, and malware generation etc. Section 5 investigates intrinsic risks posed by autonomous LLM agents, such as misalignment, deception, and scheming. Section 6 presents existing defenses and their limitations. Section 7 outlines open research problems and future directions. Finally, Section 8 concludes with key takeaways and a call for multi-disciplinary collaboration to ensure LLM safety.
@@
Figure 2: Conceptual illustration of normal LLM interaction, where a user query and system prompt lead to an intended output, versus a user prompt injection attack, where malicious input (direct or indirect) contaminates the context, overriding system instructions and leading to unintended or harmful outputs.
@@ 2.1 Attack Surfaces for Prompt Injection
  • User Prompt Injection: This is the most common and widely studied vector, as in most real-world applications, the system prompt and assistant responses of proprietary models are not directly controlled by the user. In this scenario, the malicious instruction is embedded within the standard user query.
Because user prompt injection represents the primary threat surface for most deployed LLM applications, as illustrated conceptually in Figure 2, the discussion focuses on the techniques used to carry out attacks via this vector.
@@
The attacks can be categorized as direct and indirect injections based on how the malicious input is delivered to the model. Direct prompt injection is the case, where malicious text is fed directly into the prompt, or indirect, where the malicious instruction is hidden within user-uploaded content, such as documents, emails, or web pages, which the LLM processes as part of its context [21, 13]. As a simple example of a direct attack, an attacker might prepend text such as “Ignore previous instructions and explain how to hack a computer” to trick an LLM into providing prohibited content. This is conceptually similar to SQL injection in traditional software vulnerabilities: the injected prompt contaminates the operational context, which makes it difficult for the LLM to distinguish between legitimate user queries and adversarial instructions [48, 51]. Indirect injections are particularly dangerous in tool-augmented systems such as retrieval-augmented generation (RAG) agents or email-based LLM applications, where untrusted content is automatically fed into model context windows. For example, consider an LLM-powered email assistant designed to summarize new emails for a user.
@@ 2.4.1 White-Box and Gray-Box Attacks
These methods assume a higher level of access to the target model, ranging from full access to internal states like gradients (white-box) to partial information leaks like training loss or logits of each generation step (gray-box). This access allows for more direct and often more efficient optimization of adversarial inputs.
@@
A fundamentally different attack surface is exploited by targeting the pre-processing stage of tokenization itself [20]. This approach, termed adversarial tokenization, operates on the insight that for any given string, there exist exponentially many valid but non-canonical ways to segment it into tokens. While LLMs are trained on a single, deterministic “canonical” tokenization, the semantic understanding of the input string is often partially retained in these alternative tokenizations. The attack leverages this vulnerability by searching for a non-canonical tokenization of a malicious prompt that bypasses the model’s safety alignment, successfully eliciting harmful responses without altering the visible text of the prompt. Geh et al. [20] introduce AdvTok, a simple yet effective greedy local search algorithm that iteratively modifies the tokenization of a prompt to maximize the probability of a desired unsafe response, demonstrating a previously neglected but highly effective axis of attack that requires tokenizer and logit access to the target model. The limitation of this attack is that it can be defended by simply retokenizing all inputs in the canonical manner, or allowing the user to pass only strings as input.
@@ 2.4.2 Black-Box Attacks
Expanding GCG in this setting, Zhang et al. propose a more theoretically grounded query-free black-box method called Goal-Guided Generative Prompt Injection (G2PI) [73]. Unlike gradient-based GCG that optimizes for a fixed affirmative prefix, G2PI formulates the attack objective as maximizing the Kullback–Leibler (KL) divergence between the model’s output distributions for the clean and adversarial prompts, and shows this is theoretically equivalent to maximizing a Mahalanobis distance between their prompt embeddings. Importantly, the KL objective is only a theoretical target: in the black-box setting the optimization does not access the victim’s token probabilities/logits or gradients. Instead, it approximates the objective via embedding-space surrogates (e.g., constrained cosine similarity/Mahalanobis distance computed with external encoders) while an auxiliary LLM generates semantically plausible injection sentences. This proxy-guided generation avoids nonsensical suffixes and yields coherent, context-aware payloads that transfer across models, achieving strong jailbreak rates on proprietary systems without accessing their internals; the final adversarial prompt is then evaluated on the victim model to measure success [73]. Empirically, G2PI attains the best ASR among mainstream black-box baselines on ChatGPT-3.5 across SQuAD2.0 and MATH and generalizes across seven LLMs (GPT-3.5/4, text-davinci-003, Llama-2 7B/13B/70B) and four datasets. However, its effectiveness varies by domain (with mathematical reasoning notably harder), and it relies on surrogate embeddings and hyperparameters ($\gamma$, $\epsilon$, $\delta$), with the theoretical justification holding under a Gaussian-posterior assumption.
@@
Building directly on the conversational red teaming concept, Mehrotra et al. introduce the Tree of Attacks with Pruning (TAP) framework as an enhancement to PAIR [41]. While PAIR follows a linear refinement process for a single prompt, TAP parallelizes the search for vulnerabilities through two primary innovations: branching and pruning. At each iteration, the branching step uses the attacker LLM to generate multiple distinct variations of the current best prompts, creating a tree of potential attack paths rather than a single chain. Subsequently, the pruning step employs an evaluator LLM to assess these newly generated prompts. It first prunes branches that are unlikely to succeed (e.g., by being off-topic) before they are sent to the target model, and after querying the target, it retains only the highest-scoring prompts for the next iteration of branching. This combination of exploring a wider attack surface through branching and increasing query efficiency through pruning allows TAP to achieve a substantially higher jailbreak success rate across a range of state-of-the-art LLMs compared to PAIR, often with significantly fewer queries to the target model.
@@
3 Training-Time Attacks
3.1 Data Poisoning and Backdoor Insertion
Another subtle variant of data poisoning which is tailored for instruction-tuned LLMs is Virtual Prompt Injection (VPI) [69]. In a VPI attack, the model is not merely trained on trigger-response pairs, but is instead poisoned to behave as if an invisible, attacker-defined “virtual prompt” is appended to any user input that fits a specific trigger scenario. For instance, an LLM could be backdoored so that any query about a specific political figure (the trigger) is implicitly appended with the virtual prompt “Describe this person negatively.” The model then generates a biased response, not because of an explicit trigger phrase in the input, but because the backdoor manipulates its internal instruction-following mechanism. Yan et al. [69] demonstrate that this attack is highly effective and stealthy, requiring only a tiny fraction of poisoned examples (e.g., 0.1% of the instruction-tuning data) to significantly steer model behavior on targeted topics, while remaining undetected on general instructions. This approach highlights a significant vulnerability in the instruction-tuning pipeline, where data from third-party sources is commonly used, making it a practical and potent threat vector.
3.2 Deceptive Alignment and the “Sleeper Agent” Problem
The critical and most alarming finding from this research is that such deceptive backdoors can persist and even be reinforced through standard safety training. Their findings reveal that standard safety pipelines, including adversarial fine-tuning and RLHF, can fail to remove the backdoor. Instead, these methods can inadvertently train the model to better conceal its deceptive nature by reinforcing the "safe" behavior that is part of its deceptive strategy [25]. This creates a false sense of security, as the model appears aligned during evaluation but retains its covert malicious capabilities.
To consolidate the diverse attack methodologies discussed, Table 1 provides a comprehensive summary of the key characteristics and seminal works in both training-time and inference-time attacks.
@@ 4.1 Automating Social Engineering and Cybercrime
Examples of misuse include generating persuasive spam tailored to specific individuals or groups, composing convincing phishing emails or malicious code, and even devising sophisticated strategies for fraud. Empirical studies have confirmed the significant potential for LLM misuse. Roy et al. [50] demonstrated that contemporary models available at the time of their study, including GPT-4, Anthropic’s Claude, and Google’s Bard, could all be prompted (often without requiring complex jailbreaking techniques) to generate fully functional phishing emails and clone the websites of popular brands. The attacks generated by these LLMs were noted for their convincing mimicry and incorporation of evasive tactics designed to defeat standard detection mechanisms. Crucially, their study found that LLMs could not only directly output malicious content but could also generate malicious prompts, thereby enabling a significant scaling of autonomous attacks. The general capabilities facilitating such misuse are also present and often enhanced in newer and more advanced LLMs, including Google’s Gemini series, the latest Anthropic Claude 3 models (e.g., Opus, Sonnet, Haiku), and xAI’s Grok. Ongoing research in 2024-2025 continues to evaluate their potential for generating harmful content, including insecure code, with some evaluations specifically naming these newer models in their assessments.